The points (0,3) and (3,5) lie on the graph of a linear function f. Find f(x)
Answer by stanbon(75887)

Determine the slope:
slope = m = (5-3)/(3-0) = 2/3
The form of the equation is y=mx+b.
You need to find "b"
The point (0,3) is on the line so it provides an x-value and a y-value to
plug into the form:
So substitute m=2/3, x=0, y=3 into the form to get the following:
3=(2/3)(0)+b
b= 3
The equation you want is y= (2/3)x + 3
Cheers,
Stan h